Answer the question below


In a math test, the mean score is 45 and the standard deviation is 4. Assuming normality, what is the probability that the score will lie: 

If the score is above 58

If the score is below 49

If the score is below 68

If the score is above 55

If the score is above 38


1
Expert's answer
2021-04-15T06:51:20-0400

"\mu = 45 \\

\sigma = 4 \\

P(X>58) = 1 -P(X<58) \\

= 1 -P(Z< \frac{58-45}{4}) \\

= 1 -P(Z<3.25) \\

= 1 -0.9994 \\

= 0.0006 \\

P(X<49) = P(Z< \frac{49-45}{4}) \\

= P(Z< 1) \\

= 0.8413 \\

P(X<68) = P(Z< \frac{68-45}{4}) \\

= P(Z< 5.75) \\

= 0.9999 \\

P(X>55) = 1 -P(X<55) \\

= 1 -P(Z< \frac{55-45}{4}) \\

= 1 -P(Z<2.5) \\

= 1 -0.9937 \\

= 0.0063 \\

P(X>38) = 1 -P(X<38) \\

= 1 -P(Z< \frac{38-45}{4}) \\

= 1 -P(Z< -1.75) \\

= 1 -0.04006 \\

= 0.95994"